Calculus I, Fall 2012 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 07

1. Section 4.3: problems 17, 25, 34, 75, 78, 80.

2. Section 4.4: Problems 23, 29, 35, 41, 58 (Just graph it).

3. Section 4.5: Problems 3, 5, 13, 23, 25, 26, 27(a), 33, 36.